Circular No. 5734

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

SUPERNOVA 1993I IN MCG +2-32-144 M. Turatto, C. Zanin and V. Buondi, Asiago Observatory, confirm the

existence of this object on CCD frames taken with the 1.8-m telescope on Feb. 28, when the object, located 13′′.5 west and 8′′.4 south of the galaxy’s nucleus, had B = 17.95, B − V = +0.20.

A. Clocchiatti and J. C. Wheeler, University of Texas, report that a spectrum (range 450-720 nm, dispersion 0.34 nm/pixel) obtained on Mar. 29.3 UT with the 2.7-m telescope of the McDonald Observatory shows that SN 1993I is of type Ia close to maximum light. The typical spectral features show a redshift consistent with the radial velocity of MCG +02-32-144 (12 924 km/s) reported by Zwicky in Stars and Stellar Systems 8, 375.

GRO J0422+32 C. Chevalier and S. A. Ilovaisky, Observatoire de Haute-Provence, re-

port: “New CCD photometry of the optical counterpart of this x-ray transient has been obtained with the Haute-Provence 1.2-m telescope on Jan. 19-23, Feb. 10-20 and Mar. 12-13. When combined with the earlier January data (IAUC 5692) our 1993 observations made on 23 nights over a 2.5-month interval show that, after a slow-down in its decline during early January just after the secondary x-ray maximum (IAUC 5685), the average decline rate returned to a value near 0.01 mag/day, close to what was seen earlier in the outburst (IAUC 5644, 5650). This decay rate is half that seen in V616 Mon = A0620–00 during its 1975 outburst. On 1993 Mar. 13 the optical counterpart of GRO J0422+32 was 1.4 mag fainter in V than in early Sept. 1992 when we started our monitoring program. Calibration of our differential photometry using Guide Star Photometric Catalogue stars gives V = 14.6 for this March observation. The 5.1-hr periodic modulation detected in the January data was also present in February but with a reduced amplitude.”

PERIODIC COMET SWIFT-TUTTLE (1992t) Total visual magnitude estimates: Feb. 24.76 UT, 9.2 (D. A. J. Sear-

gent, The Entrance, N.S.W., 25 × 100 binoculars); 26.76, 9.4 (P. Camilleri, Cobram, Vic., Australia, 20 × 80 binoculars); Mar. 3.76, 9.7 (Camilleri); 21.77, 10.0 (Camilleri, 0.20-m reflector); 23.74, 10.2 (Camilleri).

1993 March 30 Brian G. Marsden